OpenStudy (anonymous):

What is the solution to the equation 9^(x + 2) = 27 ?

OpenStudy (anonymous):

It is\[9^{x+2}=27\]?

hartnn (hartnn):

hint : 9=3^2 27=3^3

OpenStudy (anonymous):

@Aylin yes!

OpenStudy (anonymous):

\[9^{x+2}=9^{x} \times 9^{2}\]if that helps

OpenStudy (anonymous):

any of these @Aylin ? x = 2.5 x = 0.5 x = –0.5 x = –3.5

OpenStudy (anonymous):

I can't give you the answer, but I can help you work towards it.

hartnn (hartnn):

\(\huge 9^{x+2}= 3^{2(x+2)}=3^3\) now just equate the exponents!

hartnn (hartnn):

after equating the exponents, you get 2(x+2) =3 can you solve this for 'x' ?

OpenStudy (jiteshmeghwal9):

Because when the bases are same we can equate exponents :)

OpenStudy (hba):

Yes :)

OpenStudy (anonymous):

im really confused

hartnn (hartnn):

which step ?

OpenStudy (anonymous):

all of it

OpenStudy (jiteshmeghwal9):

First equal the bases \(\Large{{9^{(x+2)}=3^3}}\) \(\Large{(3^2)^{(x+2)}=3^3}\) gt it through here ??

OpenStudy (jiteshmeghwal9):

@Kristy143SIO ??

OpenStudy (jiteshmeghwal9):

@Kristy143SIO are u here ??

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

\(\Large{3^{2x+4}}=3^3\) now only equate the exponents 2x+4=3

OpenStudy (anonymous):

ok?

OpenStudy (jiteshmeghwal9):

solve this equation :)

OpenStudy (anonymous):